24. Use the following data to compute the real GDP for 2017 using 2010 as the base year 2010 2017 Item Price Quantity Price Quantity Catfish Pizza Taco S25 20 $12 15 20 a. $351. b. $245 c. $265 d. $197. e. $220 25. When there is a sharp increase in oil prices, the overall inflation rate is most likely rate of inflation. a. equivalent to b. below c. above d. All of the above are equally likely the core
Answer #1

24. c. $265
(Real GDP for 2017 = sum of product of quantities of 2017 and prices of 2010 of respective items = (12)*(3) + (15)*(5) + (7)*(22) = 36 + 75 + 154 = 265)

25. c. above
(overall inflation rate will exceed the core rate of inflation)